# Broken intrinsic symmetry induced magnon-magnon coupling in synthetic ferrimagnets

### v1

@@ Abstract
Synthetic antiferromagnets offer rich magnon energy spectra in which optical and acoustic magnon branches can hybridize. Here, we demonstrate a broken intrinsic symmetry induced coupling of acoustic and optical magnons in a synthetic ferrimagnet consisting of two dissimilar antiferromagnetically interacting ferromagnetic metals. Two distinct magnon modes hybridize at degeneracy points, as indicated by an avoided level-crossing. The avoided level-crossing gap depends on the interlayer exchange interaction between the magnetic layers, which can be controlled by adjusting the non-magnetic interlayer thickness. An exceptionally large avoided level crossing gap of 6 GHz is revealed, exceeding the coupling strength that is typically found in other magnonic hybrid systems based on a coupling of magnons with photons or magnons and phonons.
@@ I Introduction
Coupling between magnon modes in sAFs can lead to the exchange of energy and momentum between different magnon modes. By controlling the coupling strength, the propagation and transmission of spin waves can be manipulated. This can improve performance in magnonic devices such as magnon transistor Gong et al. (2018), and magnon logic elements Chumak (2019). The most dominant mechanisms responsible for magnon-magnon coupling are the exchange interaction (short range) MacNeill et al. (2019) and the dipolar interaction (long range) Shiota et al. (2020). In sAFs comprising two sublattices with antiferromagnetic exchange interaction, two distinct normal modes can emerge based on the relative motion of the spins precessing in the sublattices; in-phase motion (acoustic mode) or out-of-phase motion (optical mode). Generally, these modes are protected by symmetry, preventing them from coupling Patchett et al. (2022). Strategies to facilitate magnon-magnon coupling in sAFs have been studied recently Sud et al. (2020); Dai and Ma (2021); Hayashi et al. (2023); Shiota et al. (2020); Sklenar and Zhang (2021); Chen et al. (2020).
@@
Furthermore, a recent theoretical work suggested the possibility of magnon-magnon coupling for an in-plane geometry in an intrinsic symmetry-breaking mechanism using a synthetic ferrimagnetic structure Li et al. (2021). Ferrimagnets are partially compensated magnetic structures in which the sublattices with unequal magnetic moments are arranged in opposite directions, which results in non-zero spontaneous magnetization. In a synthetic structure, this can be achieved in different ways, such as by using FM layers with different saturation magnetization, $M_{\mathrm{s}}$, by using different thicknesses of the FM layers, $d$, by using an unequal number of layers, or any combination of them, all connected by negative RKKY 111also known as bilinear or quadratic interaction interaction between layers Lepadatu et al. (2017); Xie et al. (2023); Jenkins et al. (2014); Franco and Landeros (2018); Sud et al. (2023); Wang et al. (2024b); Subedi et al. (2024). Synthetic ferrimagnets can facilitate intrinsic symmetry breaking in the presence of RKKY and biquadratic interactions, which leads to the coupling of acoustic and optical modes Li et al. (2021).
@@
Figure 1: (a) Schematic illustration of the synthetic ferrimagnet comprising of a CoFe/Ru/NiFe heterostructure. The thickness of Ru is varied from 00 nm to $1$ nm to tune the coupling strength between CoFe and NiFe. (b) Illustration of the spin-torque ferromagnetic resonance (STFMR) setup, including an inset that shows the relevant torques on the magnetization $M$: $\tau_{||}$ anti-damping-like torque, $\tau_{\alpha}$ damping-like torque, $\tau_{\perp}$ field-like torque. The biasing field $H_{0}$ is applied at an angle of 45 degrees with respect to the microwave field $H_{\mathrm{rf}}$. (c) Magnetometry measurement using vibrating sample magnetometry (VSM) for a $0.8$  nm Ru thickness and the corresponding simulation results obtained using Mumax3. The simulation agrees reasonably well with the experiment. The arrows indicated the relative alignment of the two exchange-coupled ferromagnetic layers.
@@
Here, we report the experimental observation of such magnon-magnon coupling induced by broken intrinsic symmetry in synthetic ferrimagnetic structures. We use two antiferromagnetically interacting magnetic layers of similar thickness, $d$, but different saturation magnetization, $M_{\mathrm{s}}$ (CoFe and NiFe) separated by a thin spacer layer (Ru). Magnetometry using vibrating sample magnetometry (VSM) reveals the exchange interaction between the ferromagnetic layers. Using two complementary techniques, spin-torque ferromagnetic resonance (STFMR) and Brillouin light scattering (BLS), we show that the asymmetry introduced by broken intrinsic symmetry facilitates the coupling between the acoustic and optical magnon modes. Furthermore, our studies reveal that the coupling strength between the modes can be tuned by controlling the exchange interaction between the ferromagnetic layers. Our experimental results are systematically compared with micromagnetic simulations using Mumax3 Vansteenkiste et al. (2014). This study aims to leverage the magnon-magnon interaction within an in-plane geometry to develop innovative magnonic device applications.
@@ II Sample Fabrication and Experimental Details
Figure 1(a) illustrates a schematic of the synthetic ferrimagnetic sample stack. The samples comprise a trilayer of $\mathrm{Co_{90}}\mathrm{Fe_{10}(10\leavevmode\nobreak\ nm)}$/$\mathrm{Ru(}d\leavevmode\nobreak\ \mathrm{nm)}$/$\mathrm{Ni_{80}}\mathrm{Fe_{20}(10\leavevmode\nobreak\ nm)}$ with varying thickness $d$ of the non-magnetic spacer layer Ru. The structure was grown on thermally oxidized Si substrates at room temperature under an Ar pressure of $\mathrm{4.5\times{10^{-3}}}$ Torr in a magnetron sputtering system with a base pressure of $\mathrm{1.4\times{10^{-7}}}$ Torr. The Ru spacer layer was grown in a wedge shape, allowing for a continuous thickness variation from $d=0$ to 1 nm at an wedge angle of about $\mathrm{1\leavevmode\nobreak\ }\mathrm{\mu}\mathrm{deg}$, by off-center deposition to tune the interlayer exchange interaction. All samples were capped with a 5 nm $\mathrm{SiO}_{\mathrm{2}}$ layer to protect the sample from oxidation. The sample was then cut into 11 rectangular samples where the thickness of the spacer layer can be considered approximately constant over the width of each sample. The anisotropic easy axes in the FM layers were introduced along the long sides of the samples. VSM measurements were performed to characterize the static magnetic properties of the individual pieces of the sample. Figure 1(c) shows a typical VSM measurement of a synthetic ferrimagnet with negative RKKY interaction between the ferromagnetic layers, which we discuss below.
@@
To investigate the spin dynamics in the ferrimagnet, we use a combined STFMR and BLS spectroscopy approach. The STFMR signal is the spin rectification of the magnetoresistance in the device Gui et al. (2007), which provides high sensitivity measurements. In STFMR measurements a spin-polarized current in addition to an rf magnetic field exerts a torque on the magnetization, causing it to precess when the resonance condition is realized. This precession induces a time-varying magnetization, consequently giving rise to an AC resistance as a result of the anisotropic magnetoresistance or spin Hall magnetoresistance Tulapurkar et al. (2005); Sankey et al. (2006). This, combined with the AC current passed through the sample at the same frequency, generates a rectified DC voltage, which is measured using a lock-in amplifier Liu et al. (2011). For STFMR measurements, the sputtered thin films were etched to produce STFMR devices following the following steps: first, a negative-tone photoresist was used to cover 80 $\mathrm{\mu}$m $\times$ 130 $\mathrm{\mu}$m rectangular sections of the film, followed by etching using Ar ion milling of the surrounding film. The shorted coplanar waveguides were then patterned on top of the rectangular sections using optical lithography and positive photoresist. Finally, 5 nm of Ti and 20 nm of Au were e-beam evaporated and lifted off to create the waveguides on the devices as shown in Fig. 1(b).
Figure 2: Experimentally observed magnon spectra. Thermal Brillouin light scattering spectra for (a) a Ru thickness of 0.8 nm and (b) for a Ru thickness of 0.4 nm. (c) and (d) show the corresponding STFMR spectra, respectively.
@@
A bias tee was used to simultaneously apply a microwave current and measure the rectified DC voltage using a lock-in amplifier Hossain et al. (2021). A microwave frequency signal of +21 dBm power was supplied by an Agilent E8257D signal generator modulated with a frequency of 666.7 Hz. The external magnetic field $H_{\mathrm{{0}}}$ was applied in the sample plane at an angle of $45^{\circ}$ with respect to the device to achieve maximum signal strength Harder et al. (2011); Liu et al. (2011). The microwave frequency was swept from $1$ to $18$ GHz in 86 steps, with the external field being varied from $+2000$ to 00 Oe for each frequency step.
As a spin rectification measurement, STFMR probes the excited magnons indirectly. An alternative approach is thermal BLS, which directly measures the magnon population in the magnetic material at a given temperature. We used microfocused thermal BLS measurements at room temperature in a backscattering geometry using a $532$ nm continuous wave single-mode laser and a high-numerical aperture 100×100$\times$100 × objective lens ($\mathrm{NA=0.75}$). The inelastically scattered light was passed through a high-contrast multi-pass tandem Fabry Pérot interferometer with a contrast of at least $10^{15}$. The external field was swept from $+2000$ Oe to 00 Oe. For each field value, the anti-Stokes BLS spectrum was recorded. The acquisition time for the spectrum at each field was 3 min.
@@
Figure 3: Micromagnetic simulation results obtained by Mumax3. (a) Simulated spectrum for a Ru thickness of 0.8 nm. (b) Corresponding spatially-resolved simulations reveal the phase profiles of the magnon modes in the top (A) and bottom (B) layer extracted from the FFTs of the spatial simulation at (I) 40 Oe, (II) 400 Oe, and (III) 1000 Oe. Top panel (A) is the NiFe layer, bottom panel (B) is the CoFe layer. The inset shows a color wheel indicating the phase.
@@ III Results and Discussion
As is shown in Fig. 1(c), magnetometry measurements show a typical hysteresis loop of a synthetic antiferromagnet. Magnetometry measurements show secondary loops in the hysteresis for Ru thicknesses above $0.2$ nm [see Fig. S2 in the supplementary materials (SM)]. From $0.2-0.5$ nm, two secondary loops appear near the knee (the field at which the magnetization changes from saturated to unsaturated), while above 0.5 nm, a loop forms at the center [Fig. 1(c)]. The ‘knee loops’ suggest the presence of biquadratic exchange interaction between FM films for Ru thicknesses of $0.2-0.5$ nm Demokritov (1998) (see SM). The magnetometry results can be explained by carefully analyzing the expression of the volume energy density of the structure Inoue (1994):
$\mathrm{\mathscr{E}}=-\overrightarrow{\bm{\mathrm{M}}}.\overrightarrow{\bm{\mathrm{H_{0}}}}+J_{\mathrm{q}}\bm{\hat{n}}_{1}.\bm{\hat{n}}_{2}+J_{\mathrm{{bq}}}(\bm{\hat{n}}_{1}.\bm{\hat{n}}_{2})^{2}.$ (1)
Here, $\overrightarrow{\bm{M}}$ is the magnetization, $H_{\mathrm{0}}$ is the external magnetic field, $\bm{\hat{n}}_{\mathrm{i}}$ and $M_{\mathrm{Si}}$ are the magnetization unit vectors and saturation magnetization of the $i$-th layer ($i=1,2$), respectively. $J_{\mathrm{q}}$ and $J_{\mathrm{{bq}}}$ are the phenomenological energy densities for quadratic and biquadratic exchange interaction. RKKY interaction is negative (i.e., favoring antiferromagnetic alignment of the magnetic layers) when $J_{\mathrm{q}}$ is positive. At a large field, the magnetostatic energy is large enough to saturate both layers in the direction of the field [labeled ‘1’ in Fig. 1(c)]. As the field is reduced, a field is reached [labeled ‘2’ in Fig. 1(c)] where the exchange energy becomes significant and attempts to align the FM layers antiferromagnetically. As a result, the two layers reach an equilibrium at a tilted angle with respect to each other away from the saturated state, reducing the net magnetization. The field at which the magnetization starts to reduce (i.e., the knee field in the hysteresis loop) can be calculated using a macrospin model as in Eq. (1) (see SM for the derivation):
@@
where $H_{\mathrm{{0,ff}}}$ is the magnitude of the magnetic field over which the FM layers align in a forced ferromagnetic state [point ‘1’ in Fig. 1(c)].
If the biquadratic exchange interaction is negligible, magnetization changes linearly with the external field in the unsaturated region [between point ‘2’ and ‘3’ in Fig. 1(c)]] (see SM for derivation). However, strong biquadratic interaction causes a non-linear relation, creating secondary loops near the knee field in the hysteresis (see SM). As the external field decreases further, the quadratic exchange energy dominates, forming a plateau [point ‘3’ in Fig. 1(c)] with antiparallel layer alignment (see SM). The finite magnetization at this plateau results from different saturation magnetizations between the ferromagnetic layers Waring et al. (2020), creating the central loop in the magnetometry data, as seen in Fig. 1(c). The field at which the layers antiferromagnetically align is given by Eq. (2b):
$\mathrm{H_{\mathrm{{0,afm}}}=(J_{\mathrm{q}}-2J_{\mathrm{{bq}}})\left(\frac{1}{M_{\mathrm{{S2}}}}-\frac{1}{M_{\mathrm{{S1}}}}\right)},$ (2b)
with $M_{\mathrm{S1}}>M_{\mathrm{S2}}$. Equation (2b) explains the absence of a central secondary loop in magnetometry when Ru thickness is less than $0.6$ nm, given $J_{\mathrm{q}}<2J_{\mathrm{bq}}$ makes $H_{\mathrm{0,afm}}$ negative, preventing a fully antiferromagnetic state. Eqs. (2a) and (2b) provide a means to extract $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$ from magnetometry measurements when $J_{\mathrm{q}}>2J_{\mathrm{bq}}$, allowing a fully antiferromagnetic state:

### v2

@@ Abstract
Synthetic antiferromagnets offer rich magnon energy spectra in which optical and acoustic magnon branches can hybridize. Here, we demonstrate a broken intrinsic symmetry induced coupling of acoustic and optical magnons in a synthetic ferrimagnet consisting of two dissimilar antiferromagnetically interacting ferromagnetic metals. Two distinct magnon modes hybridize at degeneracy points, as indicated by an avoided level-crossing. The avoided level-crossing gap depends on the interlayer exchange interaction between the magnetic layers, which can be controlled by adjusting the non-magnetic interlayer thickness. A large avoided level crossing gap of 3.9 GHz is revealed, exceeding the coupling strength that is typically found in other magnonic hybrid systems based on a coupling of magnons with photons or magnons with phonons.
@@ I Introduction
Coupling between magnon modes in sAFs can lead to the exchange of energy and momentum between different magnon modes. By controlling the coupling strength, the propagation and transmission of spin waves can be manipulated. This can improve performance in magnonic devices such as magnon transistor Gong et al. (2018), and magnon logic elements Chumak (2019). The most dominant mechanisms responsible for magnon-magnon coupling Zhang et al. (2025) are the exchange interaction (short range) MacNeill et al. (2019) and the dipolar interaction (long range) Shiota et al. (2020). In sAFs comprising two sublattices with antiferromagnetic exchange interaction, two distinct normal modes can emerge based on the relative motion of the spins precessing in the sublattices; in-phase motion (acoustic mode) or out-of-phase motion (optical mode). Generally, these modes are protected by symmetry, preventing them from coupling Neumann and Wigner (1929); Patchett et al. (2022). In a symmetric synthetic antiferromagnet, where the two ferromagnetic layers are magnetically equivalent, a $180^{\circ}$ rotation about the field direction (when the field is aligned with a high-symmetry axis) is a valid symmetry operation. Under this operation, the acoustic mode is odd ($A\rightarrow-A$) and the optical mode is even ($A\rightarrow A$). Any coupling term must be invariant under the symmetry operation; however, the product of an even and an odd mode changes sign and therefore the coupling vanishes, allowing the modes to cross without hybridization. Strategies to facilitate magnon-magnon coupling in sAFs have been studied recently Sud et al. (2020); Dai and Ma (2021); Hayashi et al. (2023); Shiota et al. (2020); Sklenar and Zhang (2021); Chen et al. (2020).
@@
Furthermore, a recent theoretical work suggested the possibility of magnon-magnon coupling for an in-plane geometry in an intrinsic symmetry-breaking mechanism using a synthetic ferrimagnetic structure Li et al. (2021). Ferrimagnets are partially compensated magnetic structures in which the sublattices with unequal magnetic moments are arranged in opposite directions, which results in non-zero spontaneous magnetization. In a synthetic structure, this can be achieved in different ways, such as by using FM layers with different saturation magnetization, $M_{\mathrm{s}}$, by using different thicknesses of the FM layers, $d$, by using an unequal number of layers, or any combination of them, all connected by negative Ruderman–Kittel–Kasuya–Yosida (RKKY) 111also known as bilinear or quadratic interaction interaction between layers Lepadatu et al. (2017); Xie et al. (2023); Jenkins et al. (2014); Franco and Landeros (2018); Sud et al. (2023); Wang et al. (2024b); Subedi et al. (2024). Synthetic ferrimagnets can facilitate intrinsic symmetry breaking in the presence of RKKY and biquadratic interactions, which leads to the coupling of acoustic and optical modes Li et al. (2021).
@@
Figure 1: (a) Schematic illustration of the synthetic ferrimagnet comprising of a CoFe/Ru/NiFe heterostructure. The thickness of Ru is varied from 0 nm to 11 nm to tune the coupling strength between CoFe and NiFe. (b) Illustration of the spin-torque ferromagnetic resonance (STFMR) setup, including an inset that shows the relevant torques on the magnetization $\boldsymbol{M}$: $\boldsymbol{\tau}_{||}$ - anti-damping-like torque, $\boldsymbol{\tau}_{\alpha}$ - damping-like torque, $\boldsymbol{\tau}_{\perp}$ - field-like torque. The biasing field $\boldsymbol{H}_{0}$ is applied at an angle of 45 degrees with respect to the microwave field $\boldsymbol{H}_{\mathrm{rf}}$. (c) Schematic diagram of equilibrium angles of the ferromagnetic layers relative to the external field direction. (d) Vibrating sample magnetometry (VSM) measurement for a sample with Ru thickness of $0.8$ nm. Corresponding macrospin model calculations and micromagnetic simulations (MuMax3) are overlaid. Both the macrospin model and simulations show good agreement with experiment. The arrows indicate the relative alignment of the two exchange-coupled ferromagnetic layers. (e) MuMax3 simulations of the equilibrium angles of the ferromagnetic layers as a function of the applied field during field ramp-down, referenced to the external field direction defined in (c). Note that -180° and 180° correspond to the same magnetization direction. In some regions, the blue data points (magnetization direction of CoFe) are obscured by the overlapping orange data points (magnetization direction of NiFe).
@@
Here, we report the experimental observation of such magnon-magnon coupling induced by broken intrinsic symmetry in synthetic ferrimagnetic structures. We use two antiferromagnetically interacting magnetic layers of similar thickness, $d$, but different saturation magnetization, $M_{\mathrm{s}}$ (CoFe and NiFe) separated by a thin spacer layer (Ru). Magnetometry using vibrating sample magnetometry (VSM) reveals the exchange interaction between the ferromagnetic layers. Spin dynamics spectrum reveals that the asymmetry introduced by broken intrinsic symmetry facilitates the coupling between the acoustic and optical magnon modes. Furthermore, our studies show that the coupling strength between the modes can be tuned by controlling the exchange interaction between the ferromagnetic layers. Our experimental results are systematically compared with macrospin modeling and micromagnetic simulations using MuMax3 Vansteenkiste et al. (2014). Our study aims to leverage the magnon-magnon coupling within an in-plane geometry to develop innovative magnonic device applications.
@@ II Sample Fabrication and Experimental Details
Figure 1(a) illustrates a schematic of the synthetic ferrimagnetic sample stack. The samples comprise a trilayer of $\mathrm{Co_{90}}\mathrm{Fe_{10}(10\penalty 10000\ nm)}$$\mathrm{Ru(}d\penalty 10000\ \mathrm{nm)}$$\mathrm{Ni_{80}}\mathrm{Fe_{20}(10\penalty 10000\ nm)}$ with varying thickness $d$ of the non-magnetic spacer layer Ru. The structure was grown on thermally oxidized Si substrates at room temperature under an Ar pressure of $\mathrm{4.5\times{10^{-3}}}$ Torr in a magnetron sputtering system with a base pressure of $\mathrm{1.4\times{10^{-7}}}$ Torr. The Ru spacer layer was grown in a wedge shape, allowing for a continuous thickness variation from $d=0$ to 1 nm at an wedge angle of about $\mathrm{1\penalty 10000\ }\mathrm{\mu}\mathrm{deg}$, by off-center deposition to tune the interlayer exchange interaction. All samples were capped with a 5 nm $\mathrm{SiO}_{\mathrm{2}}$ layer to protect the sample from oxidation. The sample was then cut into 11 rectangular samples where the thickness of the spacer layer can be considered approximately constant over the width of each sample. The anisotropic easy axes in the FM layers were introduced along the long sides of the samples. VSM measurements were performed to characterize the static magnetic properties of the individual pieces of the sample. Figure 1(d) shows a typical VSM measurement of a synthetic ferrimagnet with negative RKKY interaction between the ferromagnetic layers, which we will discuss below.
@@
To probe the spin dynamics of the ferrimagnet, we employ spin-torque ferromagnetic resonance (STFMR) spectroscopy. The STFMR signal is the spin rectification of the magnetoresistance in the device Gui et al. (2007), which provides high sensitivity measurements. In STFMR measurements a spin-polarized current in addition to an RF magnetic field exerts a torque on the magnetization, causing it to precess when the resonance condition is realized. This precession induces a time-varying magnetization, consequently giving rise to an AC resistance as a result of the anisotropic magnetoresistance or spin Hall magnetoresistance Tulapurkar et al. (2005); Sankey et al. (2006). This, combined with the AC current passed through the sample at the same frequency, generates a rectified DC voltage, which is measured using a lock-in amplifier Liu et al. (2011). For STFMR measurements, the sputtered thin films were etched to produce STFMR devices following the following steps: first, a negative-tone photoresist was used to cover 80 $\mathrm{\mu}$m $\times$ 130 $\mathrm{\mu}$m rectangular sections of the film, followed by etching using Ar ion milling of the surrounding film. The shorted coplanar waveguides were then patterned on top of the rectangular sections using optical lithography and positive photoresist. Finally, 5 nm of Ti and 20 nm of Au were e-beam evaporated and lifted off to create the waveguides on the devices as shown in Fig. 1(b).
Figure 2: (a,b) Experimentally measured magnon spectra for Ru thicknesses of 0.8 nm and 0.4 nm, with corresponding macrospin model calculations overlaid. The color scale represents the lock-in amplifier output. (c,d) Corresponding micromagnetic (MuMax3) simulations of the STFMR spectra, with macrospin model results shown for comparison. Here, the color scale represents the FFT amplitude.
@@
A bias tee was employed to simultaneously apply a microwave current and measure the rectified DC voltage using a lock-in amplifier Hossain et al. (2021). A microwave signal with an output power of +21 dBm was generated by an Agilent E8257D signal generator and amplitude-modulated at a frequency of 666.7 Hz. The external magnetic field, $\boldsymbol{H}_{\mathrm{0}}$, was applied in the plane of the sample at an angle of $45^{\circ}$ relative to the device axis [see Fig. 1(b)] to maximize the signal amplitude Harder et al. (2011); Liu et al. (2011). For each applied field, the magnetic moments of the two ferromagnetic layers (NiFe and CoFe) reach equilibrium orientations characterized by angles $\phi_{\mathrm{a}}$ and $\phi_{\mathrm{b}}$, respectively, with respect to the external field direction $\boldsymbol{H}_{0}$ [see Fig. 1(c)]. These equilibrium angles were determined from micromagnetic simulations [see Fig. 1(e)]. The microwave frequency was swept from 1 GHz to 18 GHz in 86 steps, while the external magnetic field was varied from +200 mT to 0 mT for each frequency step.
@@
Figure 3: Micromagnetic simulations obtained with MuMax3. (a) Simulated spectrum for a Ru thickness of 0.8 nm. (b–f) Layer-resolved simulations showing the amplitude (top) and phase (bottom) profiles of the magnon modes, extracted from FFTs of the magnetization dynamics at 4, 12, 37, 56, and 100 mT. The two vertical lines in each panel indicate the resonance frequencies at that field. These lines intersect the phase spectra of CoFe (black) and NiFe (red) at the lower-frequency (solid) and higher-frequency (dashed) branches, highlighted by the corresponding horizontal markers in the phase panels. (g) Spatially resolved simulation at 12 mT with X and Y indicating the lateral dimensions of the studied devices. The color wheel represents the magnetization dynamics phase.
@@ III Results and Discussion
As is shown in Fig. 1(d), magnetometry measurements show a typical hysteresis loop of a synthetic antiferromagnet. Magnetometry measurements show secondary loops in the hysteresis for Ru thicknesses above $0.2$ nm [see Fig. S2 in the supplementary materials (SM)]. From 0.20.50.2-0.5 nm, two secondary loops appear near the knee (the field at which the magnetization changes from a saturated to an unsaturated state), while above 0.5 nm, a loop forms at the center [Fig. 1(d)]. The ‘knee loops’ suggest the presence of strong biquadratic exchange interaction between FM films for Ru thicknesses of 0.20.50.2-0.5 nm Demokritov (1998) (see SM). The magnetometry results can be explained by carefully analyzing the expression of the volume energy density of the structure Inoue (1994):
$\mathrm{\mathscr{E}}=-\boldsymbol{\mathrm{M}}\cdot\boldsymbol{\mathrm{H_{0}}}+J_{\mathrm{q}}\boldsymbol{m}_{1}\cdot\boldsymbol{m}_{2}+J_{\mathrm{{bq}}}(\boldsymbol{m}_{1}\cdot\boldsymbol{m}_{2})^{2}.$ (1)
Here, $\boldsymbol{M}$ is the magnetization, $\boldsymbol{H}_{\mathrm{0}}$ is the external magnetic field, and $\boldsymbol{m}_{\mathrm{i}}$ are the magnetization unit vectors of the $i$-th layer ($i=1,2$), respectively. $J_{\mathrm{q}}$ and $J_{\mathrm{{bq}}}$ are the phenomenological energy densities for quadratic and biquadratic exchange interaction. RKKY interaction is negative (i.e., favoring antiferromagnetic alignment of the magnetic layers) when $J_{\mathrm{q}}$ is positive. At a large field, the magnetostatic energy is large enough to saturate both layers in the direction of the field [configuration labeled ‘1’ in Fig. 1(d)]. As the field is reduced, a field is reached [beginning of configuration ‘2’ in Fig. 1(d)] where the exchange energy becomes significant and attempts to align the FM layers antiferromagnetically. As a result, the two layers reach an equilibrium at a tilted angle with respect to each other away from the saturated state, reducing the net magnetization. The critical field at which the magnetization starts to reduce (i.e., the knee field in the broader hysteresis loop) can be calculated using a macrospin model as in Eq. (1) (see SM for the derivation):
@@
where $H_{\mathrm{{0,ff}}}$ is the magnitude of the magnetic field over which the FM layers align in a forced ferromagnetic state [configuration ‘1’ in Fig. 1(d)], and $M_{\mathrm{Si}}$ are the saturation magnetizations of the $i$-th layer ($i=1,2$), respectively.
If the biquadratic exchange interaction is negligible, magnetization changes linearly with the external field in the unsaturated configuration [configuration ‘2’ in Fig. 1(d)] (see SM for derivation). However, strong biquadratic interaction causes a non-linear relation, creating secondary ‘knee loops’ near the knee field in the hysteresis (see SM).
As the external field decreases further, the quadratic exchange energy dominates, forming a plateau [configuration ‘3’ in Fig. 1(d)] with antiparallel layer alignment. The finite magnetization at this plateau results from different saturation magnetizations between the ferromagnetic layers Waring et al. (2020), creating the ‘central loop’ in the magnetometry data, as seen in Fig. 1(d). The critical field at which the layers align antiferromagnetically (i.e., the plateau region in the central part of the hysteresis loop) is given by Eq. (2b):
$H_{\mathrm{0,afm}}=(J_{\mathrm{q}}-2J_{\mathrm{bq}})\left(\frac{1}{M_{\mathrm{S2}}}-\frac{1}{M_{\mathrm{S1}}}\right),$ (2b)
where, when the external magnetic field lies between $H_{\mathrm{0,afm}}$ and $H_{\mathrm{0,afm}}$, the ferromagnetic layers align antiferromagnetically [configuration ‘3’ in Fig. 1(d)], and $M_{\mathrm{S1}}>M_{\mathrm{S2}}$. Equation (2b) explains the absence of a central loop in magnetometry when Ru thickness is less than $0.6$ nm, given $J_{\mathrm{q}}<2J_{\mathrm{bq}}$ makes $H_{\mathrm{0,afm}}$ negative, preventing a fully antiferromagnetic state. Eqs. (2a) and (2b) provide a means to extract $J_{\mathrm{q}}$ and $J_{\mathrm{bq}}$ from magnetometry measurements when $J_{\mathrm{q}}>2J_{\mathrm{bq}}$, allowing a fully antiferromagnetic state: